# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 5.A. Trees associated with an automorphism of a one-ended relatively hyperbolic group
Let $\Theta\in\phi^{n}$ with $[\mathrm{Per}(\Theta)]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ and let $h^{\prime}\in G$ loxodromic be such that $\Theta^{N}(h^{\prime})=h^{\prime}$ and that $\mathrm{Ax}(h^{\prime})$ contains $e$. As above the isometry $F_{\Theta^{N}}$ is elliptic in $T^{\phi}$. By Lemma 5.10, there exists $m\in\mathbb{N}^{*}$ such that $\Psi^{m}=\Theta^{m}$. Therefore, we see that $\mathrm{Per}(\Psi)=\mathrm{Per}(\Theta)$. Hence the conjugacy class of $H$ in $\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ is entirely determined by the edges of the fundamental domain $D$ contained in translates of axes of elements of $H$. Since $D$ is finite, and since the natural map $\mathrm{Per}_{\mathrm{NP}}(\phi^{n})\to\mathrm{Per}_{\mathrm{NP}}(\phi^{(n+1)!})$ is injective, there exist only finitely many $[H]\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ containing the conjugacy class of a loxodromic element. (In fact, their number is bounded above by the number of edges in the fundamental domain.)
@@
Let $\phi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$ and let $N\in\mathbb{N}^{*}$ be the integer given by Lemma 5.11. If $N=1$, we say that $\phi$ is almost rotationless.
Note that Lemma 5.11 implies that every element of $\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$ has an almost rotationless power.
@@ 5.B. The Periodic JSJ tree
We now need to understand vertex stabilisers of a JSJ tree given by 4.2 and its acylindricity in order to apply Theorem 2.5. We set ${\mathcal{H}}=\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$, and refer to the JSJ tree relative to ${\mathcal{P}}\cup{\mathcal{H}}$ as $T^{\mathrm{Per}}$. (Again, while this notation is neither standard nor entirely unambiguous, we use it consistently through our proofs.)
@@
Note that, by Lemma 5.11, there exists an almost rotationless power $\phi^{N}$ of $\phi$ such that every subgroup of $G$ whose conjugacy class is in ${\mathcal{H}}$ is contained in a subgroup whose conjugacy class is in $\mathrm{Per}_{\mathrm{NP}}(\Phi^{N})$. Thus, every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree is an $({\mathcal{A}},{\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\Phi^{N}))$-tree, and conversely. Thus the JSJ tree of $G$ relative to ${\mathcal{P}}\cup{\mathcal{H}}$ is also the JSJ tree relative to ${\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\Phi^{N})$ (see [GL17, Definition 2.12]). Thus, we only need to work with almost rotationless automorphisms and we will still get results regarding the periodic JSJ tree associated with an arbitrary automorphism.
Let $\phi^{N}$ be an almost rotationless power of $\phi$. Let $\mathrm{Per}_{\mathrm{NP}}(\phi^{N})=\{[H_{1}],\ldots,[H_{k}]\}$ where, for every $i\in\{1,\ldots,k\}$, the group $H_{i}$ is not elementary and there exist $\Phi_{i}\in\phi^{N}$ such that $H_{i}=\mathrm{Per}(\Phi_{i})$. Note that, for every $i\in\{1,\ldots,k\}$, 4.2(9) gives that the group $H_{i}$ fixes a unique rigid vertex $v_{i}$ in $T^{\mathrm{Per}}$ since $H_{i}$ is nonelementary.
Recall the construction of $T^{\phi}$ at the beginning of Section 5.A. Note that the trees $T^{\phi}$ and $T^{\mathrm{Per}}$ are compatible. Let $T^{\mathrm{ref}}$ be the unique minimal tree which refines $T^{\phi}$ and $T^{\mathrm{Per}}$. We denote by $p_{\phi}\colon T^{\mathrm{ref}}\to T^{\phi}$ and $p_{\operatorname{Per}}\colon T^{\mathrm{ref}}\to T^{\mathrm{Per}}$ the associated $G$-equivariant projections. The tree $T^{\mathrm{ref}}$ satisfies the following properties: a subgroup $H$ of $G$ stabilises a point in $T^{\mathrm{ref}}$ if and only if $H$ stabilises a point in both $T^{\phi}$ and $T^{\mathrm{Per}}$. Moreover, for every edge $e\in ET^{\mathrm{ref}}$, at least one of the images $p_{\phi}(e)$ or $p_{\operatorname{Per}}(e)$ is not reduced to a point.
@@
For $i\in\{1,\ldots,k\}$, let $F_{\Phi_{i}}$ be the isometry of $T^{\mathrm{ref}}$ induced by $\Phi_{i}$. As in Remark 5.6, one can show that $F_{\Phi_{i}}$ is elliptic in $T^{\mathrm{ref}}$ (this uses the acylindricity of the action of $G$ on $T^{\mathrm{ref}}$). For every $i\in\{1,\ldots,k\}$, let $T^{\mathrm{ref}}_{H_{i}}$ be the minimal tree of $H_{i}$ in $T^{\mathrm{ref}}$. It might be that every element of some $H_{i}$ is elliptic in $T^{\mathrm{ref}}$: in this case the acylindricity of $T^{\mathrm{ref}}$ implies that the whole subgroup $H_{i}$ is also elliptic. Then $H_{i}$ stabilises a vertex in both $T^{\mathrm{Per}}$ and $T^{\phi}$; since $H_{i}$ is non-elementary these vertices are unique. But since every edge of $T^{\mathrm{ref}}$ survives in the projection to at least one of $T^{\phi}$ and $T^{\mathrm{Per}}$, there cannot be an edge of $T^{\mathrm{ref}}$ stabilised by $H_{i}$, and we may take the unique fixed vertex as the minimal invariant tree in this case.
@@
Suppose that $\phi$ is almost rotationless. Let $H=\operatorname{Per}(\Phi)$, $v$ a vertex in its minimal invariant tree $T^{\mathrm{ref}}_{H}$, and $F_{\Phi}$ the isometry of $T^{\mathrm{ref}}$ induced by $\Phi$. There exists $n\in\mathbb{N}$ such that $F_{\Phi^{n}}$ fixes $v$.
@@
Let $G$ be a one-ended hyperbolic group relative to ${\mathcal{P}}$ and let $\Phi$ and $\Psi$ be two representatives of an almost rotationless $\phi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. Let $H=\operatorname{Per}(\Phi)$ and $K=\operatorname{Per}(\Psi)$ be two non-elementary periodic subgroups of $\phi$, perhaps conjugate. Then their minimal invariant trees $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ have non-empty intersection if and only if $H=K$.
@@
If $H=K$ then the minimal invariant trees $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ are equal, so only one direction needs proof. Consider the minimal invariant trees of $H$ and $K$ in $T^{\phi}$. If $T^{\phi}_{H}$ and $T^{\phi}_{K}$ do not intersect, then neither do $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$, so for the remainder of the proof we assume there is an intersection here. If the intersection contains an edge, Lemma 5.10 implies that $H=K$, so from now on assume the intersection is a single vertex $v=T^{\phi}_{H}\cap T^{\phi}_{K}$.
Each of $H$ and $K$ stabilise a unique rigid vertex in $T^{\mathrm{Per}}$, and whenever $H\neq K$ we will construct an $({\mathcal{A}},{\mathcal{P}}\cup\mathrm{Per}_{\mathrm{NP}}(\phi))$-tree where $H$ and $K$ stabilise different vertices. This prevents $\langle H,K\rangle$ being contained in a rigid vertex group, by 4.2(9). But then the rigid vertices stabilised by $H$ and $K$ are distinct, and the trees in $T^{\mathrm{ref}}$ (containing $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ as subtrees) collapsing to them must be disjoint.
We distinguish two cases, according to the nature of $v$. Note that $v$ is not QH with fibre by Corollary 5.9.
Case 1: $v$ is a rigid vertex. By Lemma 5.5 if $H$ or $K$ is elliptic and Corollary 5.8 otherwise, there exists $N\in\mathbb{N}^{*}$ such that $\operatorname{\mathrm{Stab}}(v)\subseteq\operatorname{\mathrm{Fix}}(\Phi^{N})$ and $\operatorname{\mathrm{Stab}}(v)\subseteq\operatorname{\mathrm{Fix}}(\Psi^{N})$. Hence both $\Psi^{N}$ and $\Phi^{N}$ fix elementwise the same nonelementary subgroup. By Lemma 5.1, we see that $H=K$.
Case 2: $v$ is elementary. Let ${\mathcal{T}}_{v}$ be the set of minimal trees $T_{K}^{\phi}$ with $[K]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ which contain $v$. Since, for every $[K]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, the group $K$ is nonelementary, no tree $S\in{\mathcal{T}}_{v}$ is reduced to a point. By Lemma 5.10, for all distinct $S,S^{\prime}\in{\mathcal{T}}_{v}$, the intersection $S\cap S^{\prime}$ is reduced to $v$. Let $E(v)$ be the set of edges in $S$ adjacent to $v$. We have a partition
@@
Let $S_{v}$ be the tree with one central vertex $v_{0}$ adjacent to all the other vertices and that the leaves $v_{S}$ are indexed by the trees $S\in{\mathcal{T}}_{v}$. We suppose that the stabiliser of $v_{0}$ is equal to $\operatorname{\mathrm{Stab}}(v)$ and, for every $S\in{\mathcal{T}}_{v}$, that the stabiliser of $v_{S}$ is equal to $\operatorname{\mathrm{Stab}}(E_{S})$. Let $T^{\prime}$ be the tree obtained from $T$ by blowing up $S_{v}$ at $v$ and attaching for every $e\in E$, the edge $e$ to $v_{0}$ and for every $S\in{\mathcal{T}}_{v}$ and every $e\in E_{S}$, by attaching the edge $e$ to $v_{S}$.
Note that the tree $T^{\prime}$ obtained is an $({\mathcal{A}},{\mathcal{P}})$-tree. Moreover, for any $[K_{1}],[K_{2}]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ with $T_{K_{1}}^{\phi}\neq T_{K_{2}}^{\phi}\in{\mathcal{T}}_{v}$ the minimal trees of $T_{K_{1}}^{\prime}$ and $T_{K_{2}}^{\prime}$ of $K_{1}$ and $K_{2}$ in $T^{\prime}$ are disjoint.
Let $U$ be the $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree obtained from $T^{\prime}$ by collapsing the minimal tree of every $H$ with $[H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$. Then $K$ and $H$ fix distinct points in $U$. Thus, $U$ is an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree where $H$ and $K$ fix distinct points. By Theorem 4.2 $(7)$, the groups $H$ and $K$ fix distinct rigid vertices in $T^{\mathrm{Per}}$. Therefore, the minimal trees $T_{H}^{\phi}$ and $T_{K}^{\phi}$ of $H$ and $K$ in $T^{\mathrm{ref}}$ are disjoint. ∎
@@
First note that, since the projection $p_{\operatorname{Per}}\colon T^{\mathrm{ref}}\to T^{\mathrm{Per}}$ is equivariant, the tree $T^{\mathrm{ref}}_{H}$ collapses onto $v_{H}$. Thus, we have $G_{T^{\mathrm{ref}}_{H}}\subseteq\operatorname{\mathrm{Stab}}(v_{H})$.
Conversely, let $U$ be the $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree obtained from $T^{\mathrm{ref}}$ by collapsing, for every $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, the tree $T^{\mathrm{ref}}_{H}$. For every $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, let $w_{H}$ be the vertex of $U$ fixed by $H$. Note that, by Lemma 5.15, for any $H,K\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ and every $g\in G$ such that $gHg^{-1}\neq K$, the trees $gT^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ are disjoint. Thus, the stabiliser of $w_{H}$ is equal to $G_{T^{\mathrm{ref}}_{H}}$. Since the vertex $v_{H}$ of $T^{\mathrm{Per}}$ is rigid in every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}})$-tree (see Theorem 4.2 $(7)$), the group $\operatorname{\mathrm{Stab}}(v_{H})$ is elliptic in $U$ and contains $H$. As $H$ fixes a unique point in $U$, which is $w_{H}$, we see that $\operatorname{\mathrm{Stab}}(v_{H})\subseteq G_{T^{\mathrm{ref}}_{H}}$. ∎
@@
Suppose that $\phi$ is almost rotationless. Let $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ and let $v\in VT^{\mathrm{ref}}_{H}$. Then $\operatorname{\mathrm{Stab}}(v)\subseteq H$ or the intersection $\operatorname{\mathrm{Stab}}(v)\cap H$ is elementary. In this case $\operatorname{\mathrm{Stab}}(v)$ admits a (non-reduced) splitting $\operatorname{\mathrm{Stab}}(v)=\operatorname{\mathrm{Stab}}(v)\ast_{H\cap\operatorname{\mathrm{Stab}}(v)}(H\cap\operatorname{\mathrm{Stab}}(v))$ with elementary edge stabilisers.

### 02 Mar 2026 v2

@@ 5.A. Trees associated with an automorphism of a one-ended relatively hyperbolic group
Let $\Theta\in\phi^{n}$ with $\mathrm{Per}(\Theta)]\in\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ and let $h^{\prime}\in G$ loxodromic be such that $\Theta^{N}(h^{\prime})=h^{\prime}$ and that $\mathrm{Ax}(h^{\prime})$ contains $e$. As above the isometry $F_{\Theta^{N}}$ is elliptic in $T^{\phi}$. By Lemma 5.11, there exists $m\in\mathbb{N}^{*}$ such that $\Psi^{m}=\Theta^{m}$. Therefore, we see that $\mathrm{Per}(\Psi)=\mathrm{Per}(\Theta)$. Hence the conjugacy class of $H$ in $\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ is entirely determined by the edges of the fundamental domain $D$ contained in translates of axes of elements of $H$. Since $D$ is finite, and since the natural map $\mathrm{Per}_{\mathrm{NP}}(\phi^{n})\to\mathrm{Per}_{\mathrm{NP}}(\phi^{(n+1)!})$ is injective, there exist only finitely many $H]\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$ containing the conjugacy class of a loxodromic element. (In fact, their number is bounded above by the number of edges in the fundamental domain.)
@@
Let $\phi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$ and let $N\in\mathbb{N}^{*}$ be the integer given by Lemma˜5.12. If $N=1$, we say that $\phi$ is almost rotationless.
Note that Lemma˜5.12 implies that every element of $\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$ has an almost rotationless power.
Lemma˜5.12 has also the following corollary regarding isogredience classes of automorphisms. Recall that two automorphisms $\Phi,\Psi$ of a group $G$ are in the same isogredience class if there exists $g\in G$ such that $\Phi=\mathrm{ad}_{g}\Psi\mathrm{ad}_{g}^{-1}$. The isogredience class of $\Phi$ is contained in its outer class. Lemma˜5.12 immediately implies the following.
Corollary 5.14.
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\phi=[\Phi]\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$.
There exist only finitely many isogredience classes of automorphisms in $\phi$ with a nonelementary periodic subgroup.
Notice that Lemma˜5.12 implies that there exist only finitely many isomorphism classes of nonelementary periodic subgroups for automorphisms contained in a given outer class. We ask for the following generalisation when the peripherals are virtually polycyclic.
Question 5.15.
Let $G$ be a hyperbolic group relative to virtually polycyclic groups. Do there exist only finitely many isomorphism classes of nonelementary periodic subgroups for automorphisms of $G$?
The case where the peripherals are abelian and $G$ is torsion free follows from the work of Guirardel–Levitt [28, 29].
@@ 5.B. The Periodic JSJ tree
We now need to understand vertex stabilisers of a JSJ tree given by ˜4.2 and its acylindricity in order to apply Theorem 2.5. We set ${\mathcal{H}}=\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$, and refer to the JSJ tree over elementary subgroups relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}$ as $T^{\mathrm{Per}}$. (Again, while this notation is neither standard nor entirely unambiguous, we use it consistently through our proofs.)
We highlight the fact that, if the subgroups in $\mathcal{H}$} are finitely generated, then the acylindricity and the understanding of the vertex stabilisers of $T^{\mathrm{Per}}$ mostly follow from ˜4.2 $(10)$ (using our Lemma˜5.12 for the other assumption). The main technical difficulty is thus to show that, when the periodic subgroups are not finitely generated, we still have a complete understanding of the stabilisers of rigid vertices.
@@
Note that, by Lemma 5.12, there exists an almost rotationless power $\phi^{N}$ of $\phi$ such that every subgroup of $G$ whose conjugacy class is in $\mathcal{H}$} is contained in a subgroup whose conjugacy class is in $\mathrm{Per}_{\mathrm{NP}}(\Phi^{N})$. Thus, every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}})$-tree is an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\Phi^{N}))$-tree, and conversely. Thus the JSJ tree of $G$ relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}$ is also the JSJ tree relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\Phi^{N})$ (see [30, Definition 2.12]). Therefore, we only need to work with almost rotationless automorphisms and we will still get results regarding the periodic JSJ tree associated with an arbitrary automorphism.
Let $\phi^{N}$ be an almost rotationless power of $\phi$. Let $\mathrm{Per}_{\mathrm{NP}}(\phi^{N})=\{[H_{1}],\ldots,[H_{k}]\}$ where, for every $i\in\{1,\ldots,k\}$, the group $H_{i}$ is not elementary and there exist $\Phi_{i}\in\phi^{N}$ such that $H_{i}=\mathrm{Per}(\Phi_{i})$. Note that, for every $i\in\{1,\ldots,k\}$, ˜4.2 (9) gives that the group $H_{i}$ fixes a unique rigid vertex $v_{i}$ in $T^{\mathrm{Per}}$ since $H_{i}$ is nonelementary.
Recall the construction of $T^{\phi}$ at the beginning of Section˜5.A. Note that the trees $T^{\phi}$ and $T^{\mathrm{Per}}$ are compatible. Let $T^{\mathrm{ref}}$ be the unique minimal tree which refines $T^{\phi}$ and $T^{\mathrm{Per}}$. We denote by $p_{\phi}\colon T^{\mathrm{ref}}\to T^{\phi}$ and $p_{\operatorname{Per}}\colon T^{\mathrm{ref}}\to T^{\mathrm{Per}}$ the associated $G$-equivariant projections. The tree $T^{\mathrm{ref}}$ satisfies the following properties: a subgroup $H$ of $G$ stabilises a point in $T^{\mathrm{ref}}$ if and only if $H$ stabilises a point in both $T^{\phi}$ and $T^{\mathrm{Per}}$. Moreover, for every edge $e\in ET^{\mathrm{ref}}$, at least one of the images $p_{\phi}(e)$ or $p_{\operatorname{Per}}(e)$ is not reduced to a point.
@@
For $i\in\{1,\ldots,k\}$, let $F_{\Phi_{i}}$ be the isometry of $T^{\mathrm{ref}}$ induced by $\Phi_{i}$. As in Remark 5.7, one can show that $F_{\Phi_{i}}$ is elliptic in $T^{\mathrm{ref}}$ (this uses the acylindricity of the action of $G$ on $T^{\mathrm{ref}}$). For every $i\in\{1,\ldots,k\}$, let $T^{\mathrm{ref}}_{H_{i}}$ be the minimal tree of $H_{i}$ in $T^{\mathrm{ref}}$. It might be that every element of some $H_{i}$ is elliptic in $T^{\mathrm{ref}}$: in this case Lemma˜4.3 implies that the whole subgroup $H_{i}$ is also elliptic. Then $H_{i}$ stabilises a vertex in both $T^{\mathrm{Per}}$ and $T^{\phi}$; since $H_{i}$ is non-elementary these vertices are unique. But since every edge of $T^{\mathrm{ref}}$ survives in the projection to at least one of $T^{\phi}$ and $T^{\mathrm{Per}}$, there cannot be an edge of $T^{\mathrm{ref}}$ stabilised by $H_{i}$, and we may take the unique fixed vertex as the minimal invariant tree in this case.
@@
Suppose that $\phi$ is almost rotationless. Let $\Phi\in\mathrm{NP}(\phi)$, $H=\operatorname{Per}(\Phi)$, $v$ a vertex in its minimal invariant tree $T^{\mathrm{ref}}_{H}$, and $F_{\Phi}$ the isometry of $T^{\mathrm{ref}}$ induced by $\Phi$. There exists $n\in\mathbb{N}^{*}$ such that $F_{\Phi^{n}}$ fixes v.
@@
Let $G$ be a hyperbolic group relative to $\mathcal{P}$}. Let ${\mathcal{H}}_{0}$ be a finite set of conjugacy classes of finitely generated subgroups of $G$ such that G is one-ended relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}$. Let $\Phi$ and $\Psi$ be two representatives of an almost rotationless $\phi\in\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$. Let $H=\operatorname{Per}(\Phi)$ and $K=\operatorname{Per}(\Psi)$ be two non-elementary periodic subgroups of $\phi$, perhaps conjugate. Then their minimal invariant trees $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ have non-empty intersection if and only if $H=K$.
@@
If $H=K$ then the minimal invariant trees $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ are equal, so only one direction needs proof. Consider the minimal invariant trees of $H$ and $K$ in $T^{\phi}$. If $T^{\phi}_{H}$ and $T^{\phi}_{K}$ do not intersect, then neither do $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$, so for the remainder of the proof we assume there is an intersection here. If the intersection contains an edge, Lemma˜5.11 implies that $H=K$, so from now on assume the intersection is a single vertex $v=T^{\phi}_{H}\cap T^{\phi}_{K}$.
Each of $H$ and $K$ stabilise a unique rigid vertex in $T^{\mathrm{Per}}$, and whenever $H\neq K$ we will construct an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup\mathrm{Per}_{\mathrm{NP}}(\phi))$-tree where $H$ and $K$ stabilise different vertices. This prevents $\langle H,K\rangle$ being contained in a rigid vertex group, by ˜4.2 (9). But then the rigid vertices stabilised by $H$ and $K$ are distinct, and the trees in $T^{\mathrm{ref}}$ (containing $T^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ as subtrees) collapsing to them must be disjoint.
We distinguish two cases, according to the nature of $v$. Note that $v$ is not QH with fibre by Corollary˜5.10.
Case 1: $v$ is a rigid vertex. By Lemma˜5.6 if $H$ or $K$ is elliptic and Corollary˜5.9 otherwise, there exists $N\in\mathbb{N}^{*}$ such that $G_{v}\subseteq\operatorname{\mathrm{Fix}}(\Phi^{N})$ and $G_{v}\subseteq\operatorname{\mathrm{Fix}}(\Psi^{N})$.
Hence both $\Psi^{N}$ and $\Phi^{N}$ fix elementwise the same nonelementary subgroup. By Lemma˜5.1, we see that $H=K$.
Case 2: $v$ is elementary. Let ${\mathcal{T}}_{v}$ be the set of minimal trees $T_{K}^{\phi}$ with $K]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ which contain $v$. Since, for every $K]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, the group $K$ is nonelementary, no tree $S\in{\mathcal{T}}_{v}$ is reduced to a point. By Lemma 5.11, for all distinct $S,S^{\prime}\in{\mathcal{T}}_{v}$, the intersection $S\cap S^{\prime}$ is reduced to $v$. Let $E(v)$ be the set of edges adjacent to $v$. We have a partition
@@
Let $S_{v}$ be the tree with one central vertex $v_{0}$ adjacent to all the other vertices and that the leaves $v_{S}$ are indexed by the trees $S\in{\mathcal{T}}_{v}$. We suppose that the stabiliser of $v_{0}$ is equal to $G_{v}$ and, for every $S\in{\mathcal{T}}_{v}$, that the stabiliser of $v_{S}$ is equal to the setwise stabiliser $\operatorname{\mathrm{Stab}}(E_{S})$ of $E_{S}$. Let $T^{\prime}$ be the tree obtained from $T$ by blowing up $S_{v}$ at $v$ and attaching for every $e\in E$, the edge $e$ to $v_{0}$ and for every $S\in{\mathcal{T}}_{v}$ and every $e\in E_{S}$, by attaching the edge $e$ to $v_{S}$.
Note that the tree $T^{\prime}$ obtained is an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0})$-tree. Moreover, for any $K_{1}],[K_{2}]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ with $T_{K_{1}}^{\phi}\neq T_{K_{2}}^{\phi}\in{\mathcal{T}}_{v}$ the minimal trees of $T_{K_{1}}^{\prime}$ and $T_{K_{2}}^{\prime}$ of $K_{1}$ and $K_{2}$ in $T^{\prime}$ are disjoint.
Let $U$ be the $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}})$-tree obtained from $T^{\prime}$ by collapsing the minimal tree of every $H$ with $H]\in\mathrm{Per}_{\mathrm{NP}}(\phi)$. Then $K$ and $H$ fix distinct points in $U$. Thus, $U$ is an $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}})$-tree where $H$ and $K$ fix distinct points. By Theorem 4.2 $(7)$, the groups $H$ and $K$ fix distinct rigid vertices in $T^{\mathrm{Per}}$. Therefore, the minimal trees $T_{H}^{\phi}$ and $T_{K}^{\phi}$ of $H$ and $K$ in $T^{\mathrm{ref}}$ are disjoint. ∎
@@
First note that, since the projection $p_{\operatorname{Per}}\colon T^{\mathrm{ref}}\to T^{\mathrm{Per}}$ is $G$-equivariant, the tree $T^{\mathrm{ref}}_{H}$ collapses onto $v_{H}$. Thus, we have $G_{T^{\mathrm{ref}}_{H}}\subseteq G_{v_{H}}$.
Conversely, let $U$ be the $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}})$-tree obtained from $T^{\mathrm{ref}}$ by collapsing, for every $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, the tree $T^{\mathrm{ref}}_{H}$. For every $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, let $w_{H}$ be the vertex of $U$ fixed by $H$. Note that, by Lemma 5.18, for any $H,K\in\mathrm{Per}_{\mathrm{NP}}(\phi)$ and every $g\in G$ such that $gHg^{-1}\neq K$, the trees $gT^{\mathrm{ref}}_{H}$ and $T^{\mathrm{ref}}_{K}$ are disjoint. Thus, the stabiliser of $w_{H}$ is equal to $G_{T^{\mathrm{ref}}_{H}}$. Since the vertex $v_{H}$ of $T^{\mathrm{Per}}$ is rigid in every $({\mathcal{A}},{\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}})$-tree (see Theorem 4.2 $(7)$), the group $G_{v_{H}}$ is elliptic in $U$ and contains $H$. As $H$ fixes a unique point in $U$, which is $w_{H}$, we see that $G_{v_{H}}\subseteq G_{T^{\mathrm{ref}}_{H}}$. ∎
@@
Suppose that $\phi$ is almost rotationless. For every $H\in\mathrm{Per}_{\mathrm{NP}}(\phi)$, and the unique vertex $v_{H}\in VT^{\mathrm{Per}}$ it stabilises, we have $G_{v_{H}}=H$.